OpenStudy (anonymous):

You buy a clothing at a sale. You buy sneakers at 2/3 of their original price of $60, a sweater at 7/10 of its original price of $35, and a jacket at 5/6 of its original price of $45. How much money did you spend? How much money did you save? What fraction of the total original price did you save?

OpenStudy (imsleeziboii):

nope nope nope read it wrong

OpenStudy (imsleeziboii):

the sneakers were 40 dollars

OpenStudy (imsleeziboii):

24.50$ on the sweater

OpenStudy (imsleeziboii):

64.50

OpenStudy (imsleeziboii):

and 37.50$

OpenStudy (imsleeziboii):

102$ you spent

OpenStudy (anonymous):

Can you organize it more please? @imsleeziboii

OpenStudy (shaswatrajput):

for money spend- 2/3*60 =40 7/10 *35=24.5 and 5/6 *45= 37.5 sum all these answer and get the answer for first one. For second part, add 60+35+45 and third part subtract second question answer from first and divide it with second question answer

OpenStudy (imsleeziboii):

Okay you spent $102, saved 38$

OpenStudy (imsleeziboii):

and the fraction of the origina price is 102/140

OpenStudy (imsleeziboii):

51/70 is the fraction simplified

OpenStudy (imsleeziboii):

okay so you spent 102$, saved 38$ and the fraction of all of it was 51/70

OpenStudy (imsleeziboii):

^^^

OpenStudy (anonymous):

Much better, thank you! @imsleeziboii

OpenStudy (imsleeziboii):

No problem

OpenStudy (anonymous):

Can you explain how you got it all please? @imsleeziboii

OpenStudy (imsleeziboii):

sure

OpenStudy (anonymous):

@imsleeziboii ^^

OpenStudy (imsleeziboii):

okay so \[\frac{ 2 }{ 3 }\] of 60$ is 40 dollars. \[\frac{ 7 }{ 10 }\] of 35 is 24.5 which in dollars is 24.50$ \[\frac{ 5 }{ 6 }\] of 45 is 37.5 which is 37.50$ Add all these up to get the amount spent, which is 102$ Add up all the original amounts which is 60,35,45=140 \[\frac{ 102 }{ 140 }\] simplifed is \[\frac{ 51 }{ 70 }\] because you divide each by 2 and its simplified completely

OpenStudy (imsleeziboii):

and to get how much you save you subtract how much you spent (102) from the original price of all of them (140)

OpenStudy (imsleeziboii):

okay so you spent 102$, saved 38$ and the fraction of all of it was 51/70

OpenStudy (anonymous):

Thank you! @imsleeziboii

OpenStudy (imsleeziboii):

No problem